What is the equation of the line passing through each of the given pairs of points. Result in slope-intercept form, where possible.
(-1,3) and (4,-2)

Answer by checkley75(3666)

SLOPE=(-2-3)/(4+1)=-5/5=-1 NOW REPLACE X & Y IN THE LINE FORMULA (Y=mX+b) & SOLVE FOR THE Y INTERCEPT(b):
3=-1*-1+b
3=1+b
b=3-1
b=2 FOR THE Y INTERCEPT.THUS THE LINE EQUATION IS:
Y=-X+2